OpenStudy (anonymous):

the sum of two numbers is 210. if one number is the square of the other, find the set of numbers.

OpenStudy (gudden):

let one number be x then another = x^2 and as the question says their sum is 210 so u get x^2 + x = 210 adding - 210 to both sides we get x^2 +x - 210 = 0 so try solving this equation to get the value of x

OpenStudy (mathmale):

@kedacse : Can you factor gudden's quadratic equation as a first step towars solving it for the values of x?